What is the concentration of lead ions and chromate ions in a saturated lead chromate solution at 25°C? (Ksp= 1.8 x 10-14)
Answer #1

Let S is molar concentration of Pb2+ and CrO4 2-

PbCrO4(s) -----> Pb2+(aq) + CrO4 2-(aq)

S S

Ksp = [Pb2+] [CrO4 2-]

1.8*10^-14 = S * S

S = 1.342*10^-7 M

So, [Pb2+] = [CrO4 2-] = S = 1.34*10^-7 M ....Answer